# Efficient Large-Scale Urban Parking Prediction: Graph Coarsening Based on Real-Time Parking Service Capability

### v1

@@ V Conclusion
In this study, we propose an innovative large-scale urban parking prediction framework based on real-time service capabilities. Specifically, we first introduce a GAT mechanism based on real-time service capabilities, constructing a dynamic parking graph that reflects real-time parking behavior preferences. This aids spatio-temporal GCNs models in more accurately capturing the interactions and trends of change between parking lots. Subsequently, we designed a large-scale urban parking prediction framework based on coarsening and TCN-AE, aimed at effectively reducing the training time required for models to process complex parking data. Finally, we conducted experiments on a real dataset from Shenzhen’s parking lots to validate the performance of our proposed prediction framework. The experimental results show that our framework significantly improves both accuracy and efficiency compared to traditional parking prediction methods. Furthermore, as the graph scale expands, our framework’s performance exhibits even more pronounced advantages, underscoring its immense potential in addressing complex urban parking problems in practical applications.
@@
  • Selection of Coarsening Ratio: The coarsening dimension utilized in this study was determined through experimentation with a grid search strategy, yielding a general range that may limit its applicability to the specific dataset used and not be accurate enough for others. Therefore, in our subsequent efforts, we aim to employ deep learning methods to automatically learn and ascertain the optimal coarsening dimension.
  • Integration of Multi-source Data: Given that parking demand is influenced by a wide array of factors, including nearby traffic flow, weather conditions, and more, we intend to incorporate multiple data sources for feature fusion in the future. This approach will enhance the model’s comprehension of parking scenarios, providing a richer understanding of the complex dynamics at play.
  • Global Information Aggregation: The PRGAT algorithm proposed in this paper aggregates node feature information considering only the 2nd-order neighborhood, thus overlooking the influence of distant global nodes. In our future research, we plan to incorporate higher-order global information by employing concepts from fractal theory.
@@ Acknowledgments
This work was supported by the National Key R&D Program of China (Grant No.2020YFB2104300) and Major Program of Science and Technology of Shenzhen(Grant No.KJZD20231023100304010, KJZD20231023100509018).

### v2

@@ V Conclusion
In this study, we propose an innovative large-scale urban parking prediction framework based on real-time service capabilities. Specifically, we first introduce a GAT mechanism based on real-time service capabilities, constructing a dynamic parking graph that reflects real-time parking behavior preferences. This aids spatiotemporal GCNs models in more accurately capturing the interactions and trends of change between parking lots. Subsequently, we designed a large-scale urban parking prediction framework based on coarsening and TCN-AE, aimed at effectively reducing the training time required for models to process complex parking data. Finally, we conducted experiments on a real dataset from Shenzhen’s parking lots to validate the performance of our proposed prediction framework. The experimental results show that our framework significantly improves both accuracy and efficiency compared to traditional parking prediction methods. Furthermore, as the graph scale expands, our framework’s performance exhibits even more pronounced advantages, underscoring its immense potential in addressing complex urban parking problems in practical applications.
@@
  • Cross-domain Expansion: This study currently relies on the parking dataset from Shenzhen, which is designed for the specific urban layout and traffic patterns of the city, making its applicability limited. To extend the model to other complex transportation networks (such as high-speed rail, airports, etc.), we plan to design adaptive hub node identification algorithms based on the characteristics of different domains and validate the model’s generalization ability with cross-domain datasets. This will allow the framework to better adapt to various transportation systems, improving both prediction accuracy and practical application value.
  • Memory Overhead Optimization: Currently, our approach employs a space-for-time optimization strategy, which effectively improves computational efficiency during training. However, memory consumption remains a bottleneck that needs further optimization in real-world deployment. Therefore, in future work, we plan to consider pruning the time-convolutional autoencoder by removing unnecessary layers and redundant parameters. At the same time, we will incorporate sparse regularization techniques to further compress feature representations, forcing the model to activate only the most important features, thus reducing storage requirements.
  • Global Information Aggregation: The PRGAT algorithm in this study mainly relies on feature aggregation from 2nd-order neighbors, neglecting potential connections between distant nodes. For complex transportation networks, similar functional areas in different geographical locations often exhibit similar traffic patterns at different times. Understanding these temporal and spatial pattern similarities will help improve the model’s overall understanding of traffic flow and its prediction accuracy. Therefore, in future work, we plan to introduce fractal theory to explore node behavior patterns across multiple time scales, identifying potential spatiotemporal relationships between distant nodes, thereby enhancing the model’s ability to capture spatiotemporal dependencies and improving overall prediction performance.
@@ Acknowledgments
This work was supported in part by Guangdong S&T Programme (Grant No. 2024B0101020004) and Major Program of Science and Technology of Shenzhen(Grant No.KJZD20231023100304010, KJZD20231023100509018).